OpenStudy (anonymous):

A little stuck getting started on this one...Consider the three points A = (7, -3), B = (9, 1), and C = (11, 0). Find the length of each side of the triangle ABC.

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

Do you mean C^2=a^2+b^2?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

It's the third set of points that's confusing me.

hartnn (hartnn):

length of each side means you'll have to find AB BC AC for AB, take points A and B and use the formula to get it.... AB= .... ?

OpenStudy (anonymous):

Got it! Thank you!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

and \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

I wish I found this site sooner, it has been a real big help!

hartnn (hartnn):

and it'll continue to help whenever you need :)